# Frequency- and phase-resolved polarimetry of millisecond pulsars and its application to timing

### v1

@@ 1 Introduction
Millisecond pulsars are exceptional clocks with which to carry out precision timing. The best-timed pulsars such as PSR J1909–3744 yield post-fit root-mean squared residuals $\lesssim 100\,\mathrm{ns}$. Such low-noise pulsars enable precision tests of fundamental physics (Fonseca et al., 2021; Podsiadlowski et al., 2005; Weisberg et al., 2010) and measurements of nanohertz gravitational waves (Sazhin, 1978; Detweiler, 1979), the first evidence for which has recently been put forward by several collaborations (Agazie et al., 2023; Reardon et al., 2023a; EPTA Collaboration et al., 2023; Xu et al., 2023; Miles et al., 2025).
@@
Previous multi-frequency observations have been used to investigate important concepts such as emission mechanisms, emission height in the magnetosphere and beam models mainly based on the slow pulsar population (e.g., Cordes, 1978; Rankin, 1983; Lyne & Manchester, 1988; Kijak & Gil, 2003; Johnston et al., 2008). Moreover, such studies suffer from challenges due to non-contiguous frequency coverage, the usage of different receiving systems, and data obtained at different epochs. Utilization of truly broadband instruments, such as the UWL receiver, will be invaluable to fill in the gaps, reducing systematic errors from the use of disjoint data sets. The remarkable capabilities of the UWL receiver have recently been demonstrated in wideband studies of slow pulsars dedicated to beam geometries (Jaroenjittichai et al., 2025), modelling of the polarised emission (Oswald et al., 2023b), and population statistics (Johnston et al., 2021; Oswald et al., 2023a).
@@ 2.2 Basics of pulse portraiture
with uncertainty in the $i^{\text{th}}$ frequency channel $\sigma_{\rm\psi_{i}}=\sigma_{\rm I,i}/2L_{\rm i}$. However, due to the above-mentioned noise bias, PA uncertainty in the low signal-to-noise regime (i.e., where $L_{\rm i}/\sigma_{I,i}<10$) has to be calculated numerically as described in Naghizadeh-Khouei & Clarke (1993) and Everett & Weisberg (2001).
@@ 3 Data
We pre-process the data by correcting them for RM and DM. In order to correct for DM, we generate frequency-resolved timing residuals with TEMPO2 using templates from the PPTA second data release (Curyło et al., 2023; Zic et al., 2023). We fit for DM per observation, fixing all other timing parameters. The RM in each observation is fit and corrected with the rmfit routine from PSRCHIVE. We use the brute-force method of searching for best RM, which maximises the amount of linear polarisation, and fit the RM spectrum with a Gaussian. The final set of observations is obtained by filtering data corrupted due to excessive RFI, instrumental errors and failed DM and RM fits.
@@ 4.2 Polarisation profiles, fractions and position angles
Eighteen pulsars have disjoint profile components or interpulses where the different profile components do not overlap. [^5] These include PSR J0030+0451, PSR J0614–3329, PSR J1045–4509, PSR J1125–6014, PSR J1545–4550, PSR J1603–7202, PSR J1730–2304, PSR J1744–1134, PSR J1824–2352A, PSR J1832–0836, PSR J1857+0943, PSR J1902–5105, PSR J1909–3744, PSR J1939+2134, PSR J2051–0827, PSR J2145–0750, PSR J2241–5236. Often, disjoint profile components have a high degree of linear polarisation (especially at the lowest frequencies), sometimes maintained across the whole band where they are visible. [^6] See, e.g., PSR J1545–4550, PSR J1603–7202, PSR J1730–2304, PSR J1824–2452A, and PSR J1832–0836. This is particularly interesting in light of the recent results from Kramer & Johnston (2025), who argue that radio emission can in fact be produced beyond the light cylinder along with gamma ray emission. Properties of these light-cylinder profile components are expected to be different from the ones produced in a classical way above the polar cap. A detailed study of this sample of PPTA pulsars will be presented in the follow-up paper.
The majority of PPTA pulsars have PAs with many orthogonal jumps and generally high level of variability that cannot be explained well with the rotating vector model (RVM; Radhakrishnan & Cooke, 1969). For many pulsars, the the shape of PA curves evolve with radio frequency and they are also generally flatter at phases corresponding to higher L fractions with less radio-frequency evolution (see, e.g., PSR J1024–0719 in the bottom right panel of Fig.2). What is striking is that PA curves for all pulsars are almost perfectly aligned in phase (see Fig.2). It is commonly believed that the shift in PAs is related to the height of emission (Blaskiewicz et al., 1991). Our data therefore shows that radio emission of millisecond pulsars might be generated at an extremely narrow, nearly identical height range. Further, it would mean that radius-to-frequency mapping (Cordes, 1978) which is applicable for the population of slow pulsars, does not hold in the case of millisecond pulsars.
@@
Wideband study of slow pulsars showed that pulsars with higher spin-down energy $\dot{E}$ are characterised by simpler profiles, less profile evolution, and higher fractions of L (Oswald et al., 2023a). This is supported by results from Karastergiou & Johnston (2007) who argue that these properties are related to a smaller range of emission heights, and also because magnetospheres of millisecond pulsars are much smaller due to shorter rotational periods (Kramer et al., 1998). This could result in generally more complex and evolving profile shapes. In our sample of pulsars, $\dot{E}$ covers a relatively broad range of $\log_{10}[\dot{E}/(\rm erg\,s^{-1})]~=~32~-~34$. We see that pulsars with the largest spin-down energies ($\dot{E}>2\times 10^{34}\,\rm erg\,s^{-1}$) indeed are the ones with the simplest profiles (except PSR J1824–2452A). Pulsars with the lowest $\dot{E}$ all have complex multi-component profiles. We do not see a correlation between $\dot{E}$ and polarisation fractions (Fig. 4), however this might be due to our small sample size. Moreover, median polarisation fractions across all pulsars are rather constant in the three bands (there is a slight decrease of linear polarisation and increase of absolute circular polarisation fraction). On average, PPTA pulsars are linearly polarised in $\sim$20% and circularly polarised in $\sim 5$%, which is consistent with the population of slow pulsars (Oswald et al., 2023a; Posselt et al., 2023).
@@ 5 Summary and conclusions
    Comparing TOAs obtained with our new polarimetric templates and total intensity templates, we see that the improvement in TOA uncertainties is up to $\sim$20-30%. Considering the fact that currently root-mean-square of the residuals is between several $\mu s$ to hundreds of ns (e.g., Reardon et al., 2023b) and the deviations induced by gravitational waves are of the order of tens of ns, reducing the measurement uncertainties even by $\sim$100-200 ns is extremely valuable. As a large fraction of PPTA pulsars is bright, has high levels of polarisation and profile evolution in all Stokes parameters, we expect polarimetric timing to be highly beneficial for the majority of them.
@@ Acknowledgements
Murriyang, CSIRO’s Parkes radio telescope, is part of the Australia Telescope National Facility (https://ror.org/05qajvd42) which is funded by the Australian Government for operation as a National Facility managed by CSIRO. We acknowledge the Wiradjuri people as the Traditional Owners of the Observatory site. Parts of this work was completed with support from the Australian Research Council (ARC) Centre of Excellence CE230100016. MEL is supported by an Australian Research Council (ARC) Discovery Early Career Research Award DE250100508. This work was performed on the OzSTAR national facility at Swinburne University of Technology. The OzSTAR program receives funding in part from the Astronomy National Collaborative Research Infrastructure Strategy (NCRIS) allocation provided by the Australian Government.

### v2

@@ 1 Introduction
Millisecond pulsars are exceptional clocks with which to carry out precision timing. The best-timed pulsars such as PSR J1909–3744 yield post-fit root-mean squared residuals $\lesssim 100\,\mathrm{ns}$. Such low-noise pulsars enable precision tests of fundamental physics (Fonseca et al., 2021; Podsiadlowski et al., 2005; Weisberg et al., 2010) and measurements of nanohertz gravitational waves (Sazhin, 1978; Detweiler, 1979), the first evidence for which has recently been put forward by several collaborations (Agazie et al., 2023; Reardon et al., 2023b; EPTA Collaboration et al., 2023; Xu et al., 2023; Miles et al., 2025).
@@
Previous multi-frequency observations have been used to investigate important concepts such as emission mechanisms, emission height in the magnetosphere and beam models mainly based on the slow pulsar population (e.g., Cordes, 1978; Rankin, 1983; Lyne and Manchester, 1988; Kijak and Gil, 2003; Johnston et al., 2008). Moreover, such studies suffer from challenges due to non-contiguous frequency coverage, the usage of different receiving systems, and data obtained at different epochs. Utilization of truly broadband instruments, such as the UWL receiver, will be invaluable to fill in the gaps, reducing systematic errors from the use of disjoint data sets. The remarkable capabilities of the UWL receiver have recently been demonstrated in wideband studies of slow pulsars dedicated to beam geometries (Jaroenjittichai et al., 2025), modelling of the polarised emission (Oswald et al., 2023b), and population statistics (Johnston et al., 2021; Oswald et al., 2023a).
@@ 2.2 Basics of pulse portraiture
with uncertainty in the $i^{\text{th}}$ frequency channel $\sigma_{\rm\psi_{i}}=\sigma_{\rm I,i}/2L_{\rm i}$. However, due to the above-mentioned noise bias, PA uncertainty in the low signal-to-noise regime (i.e., where $L_{\rm i}/\sigma_{I,i}<10$) has to be calculated numerically as described in Naghizadeh-Khouei and Clarke (1993) and Everett and Weisberg (2001).
@@ 3 Data
We pre-process the data by correcting them for RM and DM. In order to correct for DM, we generate frequency-resolved timing residuals with TEMPO2 using templates from the PPTA third data release (Curyło et al., 2023; Zic et al., 2023). We fit for DM per observation, fixing all other timing parameters. The RM in each observation is fit and corrected with the rmfit routine from PSRCHIVE. We use the brute-force method of searching for best RM, which maximises the amount of linear polarisation, and fit the RM spectrum with a Gaussian. The final set of observations is obtained by filtering data corrupted due to excessive RFI, instrumental errors and failed DM and RM fits.
@@ 4.2 Polarisation profiles, fractions and position angles
Eighteen pulsars have disjoint profile components or interpulses where the different profile components do not overlap. [^5] These include PSR J0030+0451, PSR J0614–3329, PSR J1045–4509, PSR J1125–6014, PSR J1545–4550, PSR J1603–7202, PSR J1730–2304, PSR J1744–1134, PSR J1824–2352A, PSR J1832–0836, PSR J1857+0943, PSR J1902–5105, PSR J1909–3744, PSR J1939+2134, PSR J2051–0827, PSR J2145–0750, PSR J2241–5236. Often, disjoint profile components have a high degree of linear polarisation (especially at the lowest frequencies), sometimes maintained across the whole band where they are visible. [^6] See, e.g., PSR J1545–4550, PSR J1603–7202, PSR J1730–2304, PSR J1824–2452A, and PSR J1832–0836. This is particularly interesting in light of the recent results from Kramer and Johnston (2025), who argue that radio emission can in fact be produced beyond the light cylinder along with gamma ray emission. Properties of these light-cylinder profile components are expected to be different from the ones produced in a classical way above the polar cap. A detailed study of this sample of PPTA pulsars will be presented in the follow-up paper.
The majority of PPTA pulsars have PAs with many orthogonal jumps and generally high level of variability that cannot be explained well with the rotating vector model (RVM; Radhakrishnan and Cooke, 1969). For many pulsars, the the shape of PA curves evolve with radio frequency and they are also generally flatter at phases corresponding to higher L fractions with less radio-frequency evolution (see, e.g., PSR J1024–0719 in the bottom right panel of Fig.2). What is striking is that PA curves for all pulsars are almost perfectly aligned in phase (see Fig.2). It is commonly believed that the shift in PAs is related to the height of emission (Blaskiewicz et al., 1991). Our data therefore shows that radio emission of millisecond pulsars might be generated at an extremely narrow, nearly identical height range. Further, it would mean that radius-to-frequency mapping (Cordes, 1978) which is applicable for the population of slow pulsars, does not hold in the case of millisecond pulsars.
@@
Wideband study of slow pulsars showed that pulsars with higher spin-down energy $\dot{E}$ are characterised by simpler profiles, less profile evolution, and higher fractions of L (Oswald et al., 2023a). This is supported by results from Karastergiou and Johnston (2007) who argue that these properties are related to a smaller range of emission heights, and also because magnetospheres of millisecond pulsars are much smaller due to shorter rotational periods (Kramer et al., 1998). This could result in generally more complex and evolving profile shapes. In our sample of pulsars, $\dot{E}$ covers a relatively broad range of $\log_{10}[\dot{E}/(\rm erg\,s^{-1})]~=~32~-~34$. We see that pulsars with the largest spin-down energies ($\dot{E}>2\times 10^{34}\,\rm erg\,s^{-1}$) indeed are the ones with the simplest profiles (except PSR J1824–2452A). Pulsars with the lowest $\dot{E}$ all have complex multi-component profiles. We do not see a correlation between $\dot{E}$ and polarisation fractions (Fig. 4), however this might be due to our small sample size. Moreover, median polarisation fractions across all pulsars are rather constant in the three bands (there is a slight decrease of linear polarisation and increase of absolute circular polarisation fraction). On average, PPTA pulsars are linearly polarised in $\sim$20% and circularly polarised in $\sim 5$%, which is consistent with the population of slow pulsars (Oswald et al., 2023a; Posselt et al., 2023).
@@ 5 Summary and conclusions
    Comparing TOAs obtained with our new polarimetric templates and total intensity templates, we see that the improvement in TOA uncertainties is up to $\sim$20-30%. Considering the fact that currently root-mean-square of the residuals is between several $\mu s$ to hundreds of ns (e.g., Reardon et al., 2023a) and the deviations induced by gravitational waves are of the order of tens of ns, reducing the measurement uncertainties even by $\sim$100-200 ns is extremely valuable. As a large fraction of PPTA pulsars is bright, has high levels of polarisation and profile evolution in all Stokes parameters, we expect polarimetric timing to be highly beneficial for the majority of them.
@@ Acknowledgements
We thank Simon Johnston for valuable discussions and feedback. Murriyang, CSIRO’s Parkes radio telescope, is part of the Australia Telescope National Facility (https://ror.org/05qajvd42) which is funded by the Australian Government for operation as a National Facility managed by CSIRO. We acknowledge the Wiradjuri people as the Traditional Owners of the Observatory site. Parts of this work was completed with support from the Australian Research Council (ARC) Centre of Excellence CE230100016. MEL is supported by an Australian Research Council (ARC) Discovery Early Career Research Award DE250100508. This work was performed on the OzSTAR national facility at Swinburne University of Technology. The OzSTAR program receives funding in part from the Astronomy National Collaborative Research Infrastructure Strategy (NCRIS) allocation provided by the Australian Government.